Question 2) Transform the given wye network to equivalent delta network. 12022 10022 15022
Answer #1

RA M mi 120 M 100 R 150 R R₂ RA R1 + R2+ R3 (120x1001 3001 * RA = 120+ 100+ 150 R2 R3 RB = R2 + R3t RY (loox 150 37502 → RB 1